### v2

# Deep Instruction Tuning for Segment Anything Model

Abstract.
Recently, Segment Anything Model (SAM) has become a research hotspot in the fields of multimedia and computer vision, which exhibits powerful yet versatile capabilities on various (un) conditional image segmentation tasks. Although SAM can support different types of segmentation prompts, we note that, compared to point- and box-guided segmentations, it performs much worse on text-instructed tasks, e.g., referring image segmentation (RIS). In this paper, we argue that deep text instruction tuning is key to mitigate such shortcoming caused by the shallow fusion scheme in its default light-weight mask decoder. To address this issue, we propose two simple yet effective deep instruction tuning (DIT) methods for SAM, one is end-to-end and the other is layer-wise. With minimal modifications, DITs can directly transform the image encoder of SAM as a stand-alone vision-language learner in contrast to building another deep fusion branch, maximizing the benefit of its superior segmentation capability. Extensive experiments on three highly competitive benchmark datasets of RIS show that a simple end-to-end DIT can improve SAM by a large margin, while the layer-wise DIT can further boost the performance to state-of-the-art with much less data and training expenditures. Our code is released at: https://github.com/wysnzzzz/DIT.
footnotetext: †Corresponding Author.
1. Introduction
Figure 1. Illustration of the default fusion of SAM (a) and our deep instruction tuning (DIT) methods (b), i.e., the end-to-end (red) and layer-wise (blue) ones. Compared with the default shallow fusion, DITs regard the visual encoder as an deep multi-modal learner for cross-modal fusion and visual feature learning, thereby achieving the full interactions between text and image for text-guided SAM.
Motivated by the great success of large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023), Kirillov et al. (Kirillov et al., 2023) recently propose a Segment Anything project to build a foundation model for image segmentation, termed Segment Anything Model (SAM). SAM demonstrates an impressive yet versatile segmentation capacity by pre-training on over one billion masks. Moreover, it can support interactive segmentation conditioned on various input prompts (Kirillov et al., 2023), such as points, boxes or texts. Its outstanding ability also arouses great interest from the communities of multimedia (Cheng et al., 2023a, b; Ye et al., 2023; Gao et al., 2023) and computer vision (Wu et al., 2023a; Deng et al., 2024; Cheng et al., 2024), and has been recently transferred to various segmentation tasks, such as medical image segmentation (Guo et al., 2024; Xie et al., 2024; Quan et al., 2024), urban image segmentation task (Zhang et al., 2024a), and marine animal segmentation (Zhang et al., 2024b), or to provide instance masks for image editing tasks (Xie et al., 2023; Gao et al., 2023).
However, the text-awareness of SAM is much inferior than following the prompts of points or boxes (Kirillov et al., 2023). For instance, on true-color segmentation task (Xiao et al., 2024), SAM can achieve state-of-the-art performance of 85.2% mIoU with box guidance. In stark contrast, SAM can only achieve 58.0% mIoU when fine-tuned on the RefCOCO benchmark (Yu et al., 2016) of the popular text instructed task, i.e., referring image segmentation (RIS) (Yu et al., 2016) as shown in Fig. 2. To explain, text instructions naturally exhibit linguistic ambiguities, while location prompts (points and boxes) are precise. Meanwhile, recent Transformer-based vision models (Dosovitskiy et al., 2020; Liu et al., 2021b) have well incorporated with this type of spatial information, e.g., position embedding, and it is natural for SAM to well master these location prompts. In addition, the collection of instance-wise text descriptions is more laborious, and even the public Segment Anything project also does not release the text-conditioned annotations (Kirillov et al., 2023).
Even worse, the shallow fusion scheme in SAM further aggravates such weakness and degrades its multi-modal ability. Concretely, for either point, box or text, input prompt is first encoded by the prompt encoder of SAM, then it is directly fed into the light-weight encoder for mask prediction. In SAM, prompts can only interact to a limited extent with visual tokens via two cross-attention layers in its mask decoder (Kirillov et al., 2023). Although such mechanism is well suited for point and box prompts which are relatively easy to follow, it is insufficient for SAM to infer the intentions behind texts step-by-step. Oftentimes, SAM needs to resolve linguistic ambiguity first through the attributes and relationships of the referent, then the object can be segmented using its powerful and class-agnostic segmentation ability. To this end, we argue that
“Deep text instruction tuning is essential for SAM.”
To validate this argument, we propose two deep instruction tuning (DIT) methods for SAM. Different from existing RIS models  (Huang et al., 2020; Deng et al., 2021; Kim et al., 2022; Li et al., 2022) that build deep fusion branches upon image encoders, we focus on improving text instruction following ability of SAM without modifying its structure. This property can help DITs make full use of its superior segmentation capability while reducing the training cost to a large extent. In particular, we directly extend SAM to a deep vision-language model. Enabled by the flexibility of Transformer-based encoder (Dosovitskiy et al., 2020), we convert text instruction words to input tokens and concatenate them with the visual ones for deep multi-modal fusion and feature learning, which forms the first DIT scheme in this paper, we termed end-to-end DIT (E-DIT) , as depicted in Fig. 3. This simple method improves SAM by a large margin, e.g., +11.8% on RefCOCO val, as shown in Fig. 2. To further enhance instruction learning, we also propose a layer-wise instruction tuning method (L-DIT) that independently projects and inserts the text words into each layer of SAM. This layer-wise method has a similar principle with soft prompt tuning (Jia et al., 2022). The difference is that the inserted tokens are the projected text embeddings that vary across examples, instead of task-specific and parameterized prompt tokens that are static. Besides, DITs aim to project the text words onto the well-learned visual semantic space of SAM, while prompt tuning is often to help the model better adapt the downstream task without full tuning (Jia et al., 2022).
Figure 2. The comparison among the default fusion scheme of SAM (with ViT-B) and our deep instruction tuning methods on RefCOCO (Yu et al., 2016), i.e., the end-to-end and layer-wise ones, respectively [^2] Three methods are trained using the frozen ViT-B encoder.. In SAM, the text prompts only interact with the visual tokens in the shallow mask decoder, which is insufficient for cross-modal fusion and text-guided segmentation. Our simple deep tuning methods can improve the performance by a large margin without significantly changing the architecture of SAM.
With the proposed DIT methods, we apply SAM to the popular text-guided Referring Image Segmentation (RIS) (Liu et al., 2023c) task, and term the new models as DIT-SAMs. Extensive experiments are conducted on three highly competitive RIS benchmarks, namely RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016). The experimental results not only show the great advantages of DIT-SAMs over the default SAM, e.g., up to 11.6% improvement on RefCOCO testA, but also confirm the competitiveness of DIT-SAM as a stand alone RIS model. Overall, we provide a fast and feasible solution to extend SAM’s modal-modal ability.
To summarize, our contributions are three-fold:
  • We reveal the key drawbacks of SAM in terms of text-guided segmentation and argue that deep text instruction tuning is key to success.
  • We propose two simple yet effective deep instruction tuning (DIT) methods for SAM, which are end-to-end and layer-wise, respectively. These DIT methods pave a quick and feasible way for the multi-modal extension of SAM.
  • The proposed DIT-SAMs not only greatly outperforms the default SAM on RIS benchmarks, but also yield strong competitiveness against existing SOTA methods of RIS.
2. Related Work
2.1. Image Segmentation
Image segmentation (Ronneberger et al., 2015; Long et al., 2015; He et al., 2017; Badrinarayanan et al., 2017) is a fundamental task in multimedia and computer vision, which requires pixel-level understanding of a given image. Recent years have witnessed its fast development supported by a bunch of segmentation-related tasks, such as semantic segmentation (Long et al., 2015; Badrinarayanan et al., 2017; Chen et al., 2017; He et al., 2017; Xie et al., 2021) which categorizes individual pixels into a predetermined set of classes, instance segmentation which primarily concerns the recognition and delineation of distinct object instances (He et al., 2017; Tian et al., 2020; Wang et al., 2020b), and panoptic segmentation which merges semantic and instance segmentation tasks by assigning class labels and instance identification (Kirillov et al., 2019; Li et al., 2019). In addition, a set of advanced segmentation networks (He et al., 2017; Tian et al., 2020; Wang et al., 2020b; Ma et al., 2021; Zu et al., 2023) are also proposed to promote the development of image segmentation. For instance, He et al. (He et al., 2017) propose Mask-RCNN for object instance segmentation. SOLO (Wang et al., 2020a) converts instance segmentation into a single-shot classification-solvable problem. IFR (Ma et al., 2021) is a implicit feature refinement module for high-quality instance segmentation, and TAR (Zu et al., 2023) presents the first attempt of weakly-supervised text instance segmentation by text recognition and segmentation. Recently, the Segment Anything project (Kirillov et al., 2023) proposes a powerful and general model for image segmentation, termed segment anything model (SAM). This project includes the creation of an unprecedentedly large segmentation dataset, so that SAM is trained with about 1 billion image-mask pairs. Via large-scale pre-training, SAM exhibits robust and versatile performance across various segmentation challenges which can also handle different types of prompts for conditional segmentation, such as point, box and text. More recently, SAM has been widely introduced to a set of segmentation-related tasks, such as low-level structural segmentation (Chen et al., 2023), surgical scene segmentation (Paranjape et al., 2023), and medical image segmentation (Guo et al., 2024; Xie et al., 2024; Quan et al., 2024). However, these applications mainly make use of the interactive capability of SAM on point or box prompts, while the text-instructed application is barely explored. Moreover, SAM also works much worse with texts than the other types of prompts due to the complexity of text information and the insufficient fusion of its default shallow decoder. In this paper, we propose deep instruction tuning for the multi-modal enhancement of SAM.
2.2. Referring Image Segmentation
Referring image segmentation (RIS) (Yu et al., 2016) is one of the most popular text-conditioned segmentation task (Luo et al., 2020b, a; Jiao et al., 2021; Zhu et al., 2022; Shi et al., 2023; Liu et al., 2023f), which grounds the referent with binary masks according to the given natural language expression. Early research in RIS (Yu et al., 2016; Liu et al., 2019) typically adopts a two-stage pipeline and frames RIS as a region-text matching problem. One-stage RIS models (Liu et al., 2017; Li et al., 2018; Margffoy-Tuay et al., 2018) recently garner growing attention from both academia and industry. In these one-stage models, text features are often embedded into the vision network and directly output the mask of the referent. Inspired by the success of Transformer, various Transformer-based RIS models (Zhu et al., 2022; Hu et al., 2023; Liu et al., 2023d, a; Wu et al., 2023c; Xu et al., 2023b) are recently proposed to improve the multi-modal reasoning ability for RIS. VLT (Ding et al., 2022) and SeqTR (Zhu et al., 2022) use a encoder-decoder framework to fuse visual and language inputs into the decoder. ReSTR (Kim et al., 2022) employs two separate Transformer encoders to handle two different modalities. More recently, attempts have also been made to explore a wider range of text conditional segmentation by combining large language models (LLMs) and training with large-scale image-text pairs (Liu et al., 2023g; Yan et al., 2023; Lai et al., 2023; Zou et al., 2023). For instance, UniLSeg (Liu et al., 2023g), LISA (Lai et al., 2023) and SEEM (Zou et al., 2023) focus on text-instructed open-world segmentation, which typically require massive multi-modal training data and LLMs like LLaMA-7B (Touvron et al., 2023). Compared with these advancements, we mainly focus on addressing the text awareness of SAM and only conduct quick validations on RIS benchmarks. Among these progresses, the most relevant works to our DITs are LAVT (Yang et al., 2022) and ETRIS (Li and Sigal, 2021), which also only use two modality encoders for deep multi-modal fusion. However, our DITs greatly differ in both principle and designs. In particular, LAVT and ETRIS aim to achieve multi-modal interactions via cross-attention-based modules between two Transformers. In contrast, the principle of DITs is to project the text words onto the semantic space of SAM’s image encoder, thus achieving feature fusion and learning in one forward network. This property also makes DIT’s designs more intuitve and simpler, which also exhibits obvious merits in both training costs and performance.
Figure 3. Overall pipeline of Deep Instruction Tuning (DIT). We propose (a) End-to-end DIT (E-DIT) that appends language instructions before visual features to achieve early-fusion for enhancing multi-modal reasoning capability; (b) Layer-wise DIT (L-DIT) that layer-wisely projects to text tokens interact with the existing sequences in the well learned semantic space.
2.3. Prompt Tuning
With the ever increasing parameter size of pre-trained models (Devlin et al., 2018; Dosovitskiy et al., 2020; Touvron et al., 2023), the direct fine-tune on downstream tasks becomes prohibitively expensive. In this case, prompt tuning (Gu et al., 2021; Liu et al., 2021a; Wei et al., 2021; Han et al., 2022) is proposed to use hand-craft or learnable prompt tokens to mitigate distribution shifts between pre-training and downstream tasks. The significant achievements of prompt tuning in natural language processing (NLP) have also inspired its application to computer vision and vision-language research (Zhou et al., 2022b, a; Wu et al., 2023b; Luo et al., 2023; Liu et al., 2023e). CoOp (Zhou et al., 2022b) remains CLIP parameters unchanged and uses learnable vectors as soft prompts for the input text, thereby maximizing CLIP’s zero-shot retrieval capabilities. To address the overfitting issue on base classes, CoCoOp (Zhou et al., 2022a) proposes instance adaptation by incorporating the visual features of each image into the parameterized prompts. RPO (Lee et al., 2023) enhances the robustness of VL models by masked attention to prevent the added learnable embeddings from altering the original internal representations of the model. More recently, Wu et al. (Wu et al., 2023b) propose a method of approximated prompt tuning to reduce the additional cost of the inserted prompt tokens. Regarding soft prompt tuning (Lester et al., 2021; Luo et al., 2023; Wu et al., 2023d), whether applied deeply or shallowly, it employs learnable tokens to assist the model in adapting to downstream data distributions without the necessity for complete tuning. Although the proposed DIT methods and prompt tuning are quite similar in process, but they still hold different principles and designs. To explain, prompt tuning often adopts learnable or hand-craft tokens to brdige the gap between pre-training and downstream tasks, which are often static to tasks (Jia et al., 2022). In contrast, DITs aim to project the text tokens onto the semantic space of SAM, which are dynamic for different examples. Conclusively, prompt tuning primarily facilitates rapid task adaptation, whereas our DITs offer a novel approach to VL reasoning for SAM.
3. Method
3.1. Problem Definition
Referring image segmentation (RIS) (Yu et al., 2016) aims to segment the referred object in an image $I$ conditioned on the paired natural language expression $T$, its objective is formulated as
(1) $\arg\min_{\theta}\mathcal{L}(\mathcal{F}(I,T;\theta),M),$
where $\mathcal{F}$ denotes the model with its parameters $\theta$, $M\in\mathbb{R}^{H\times W}$ is the ground-truth mask, and $\mathcal{L}$ is the canonical cross entropy loss.
To solve this task, SAM (Kirillov et al., 2023) first extracts visual and text features through its vision and language encoders, based on which the light-weight decoder takes as input to predict the binary masks. Thus, the model $\mathcal{F}$ instantiated by SAM is
(2) $\mathcal{F}(I,T;\theta)=\mathcal{F}_{d}\big{(}\mathcal{F}_{v}(I),\mathcal{F}_{t}(T)\big{)},$
where $\mathcal{F}_{v}$, $\mathcal{F}_{t}$ and $\mathcal{F}_{d}$ refer to visual backbone, language encoder and mask decoder, respectively.
The mask decoder is very lightweight and only contains two cross-attention layers, which are in charge of both mask decoding and multi-modal interactions. In this case, SAM can be regarded as one of the late-fusion models which proved to be less effective than the early-fusion ones (Luo et al., 2020a, b; Zhu et al., 2022; Liu et al., 2023c) in addressing the complicated multi-modal reasoning tasks. Meanwhile, due to the lack of training on massive cross-modal data, text-guided SAM performs worse than the point- or box-guided SAM, leading to much inferior cross-modal abilities than existing SOTA methods in RIS (Zhu et al., 2022; Liu et al., 2023c).
We introduce Deep Instruction Tuning (DIT) for SAM to better follow the intentions reflected in the natural language expression $T$. In particular, DIT keeps the whole SAM architecture unchanged and directly uses the image encoder as a strong multi-modal fusion network. The text instruction words can continually interact with the visual contents through the whole feature learning process, helping SAM progressively locate the referent. Specifically, SAM with DIT can be then formulated as
(3) $\mathcal{F}(T,I;\theta)=\mathcal{F}_{d}\big{(}\mathcal{F}_{v}(I,\phi(\mathcal{F}_{t}(T))),\mathcal{F}_{t}(T)\big{)}.$
Here, $\phi(\cdot)$ is a mapping function to project the text features onto the default visual space of SAM.
3.2. End-to-end Deep Instruction Tuning
One straightforward solution for SAM is to directly feed the instruction words to its ViT-based image encoder. Due to the flexibility of Transformer architecture, the text features can be also regarded as additional input tokens for feature transformations. This attempt is also validated in previous VL works like ViLT (Kim et al., 2021), and has a similar process with the popular prompt tuning (Jia et al., 2022).
Here, we term this DIT solution as the end-to-end tuning, abbreviated as E-DIT. Concretely, given a text instruction $T$, we first use the language encoder, i.e., BERT (Devlin et al., 2018), to extract its representations, denoted as $\mathbf{F}_{t}\in\mathbb{R}^{l\times d}$, where $l$ denotes the length of instruction. The visual features after the patch embedding layer are denoted as $\textbf{F}_{v}^{0}\in\mathbb{R}^{m\times d}$. Here, $m=H\times W/P^{2}$ is the number of patches, where $H\times W$ is the image resolution and $P^{2}$ is the patch size. Afterwards, we use a linear layer to project $\mathbf{F}_{t}$ onto the visual semantic space of SAM, and obtain $\mathbf{F}_{t}^{0}$ via
(4) $\mathbf{F}_{t}^{0}=\mathbf{F}_{t}\textbf{W}_{0}+b_{0},$
where $\textbf{W}_{0}\in\mathbb{R}^{d\times d}$ and $b_{0}\in\mathbb{R}^{d}$ denote the projection weight and the bias term, respectively.
These text tokens and image patches are then processed by the following Transformer layers of the encoder for both image feature learning and cross-modal fusion. The output image features $\mathbf{F}_{v}^{n}$ alone with the text features $\mathbf{F}_{t}$ are fed to the lightweight mask decoder. In this case, E-DIT is formulated by
(5) $\mathbf{F}_{m}^{0}=[\mathbf{F}_{t}^{0};\mathbf{F}_{v}^{0}],$
(6) $\mathbf{F}_{m}^{i+1}=L_{i}(\mathbf{F}_{m}^{i}),$
(7) $M^{\prime}=\mathcal{F}_{d}(\mathbf{F}_{v}^{n},\mathbf{F}_{t}\mathbf{W}_{o}).$
Here, $L_{i}$ denotes the i-th Transformer layer of visual backbone, and $W_{o}\in\mathbf{R}^{d\times d}$ is the projection weight. []delimited-[][$\cdot$][ ⋅ ] denotes concatenation. $M^{\prime}$ is the predicted binary mask.
Considering the case of freezing SAM’s backbone, Eq. 4 only adds a few parameters for semantic projection. When the hidden Transformer layers are frozen, the inserted text tokens do not fit well into visual space. For the output of each layer, we also add a linear projection to facilitate text adaption.
Table 1. Comparison of different text-instructed schemes on RefCOCO. Default refers to the default lightweight fusion scheme of SAM. Add dec layers is to increasing the depth of the mask decoder. * denotes that the image backbone is frozen, while the text and mask decoders are tuned. The best and second best performances are in bold and underlined.
Method val testA testB
P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU
Default 72.20 59.56 18.78 60.74 63.41 75.28 63.76 18.76 62.95 65.65 66.95 52.63 19.82 56.69 59.95
Default 64.77 51.92 16.82 55.67 57.98 68.29 56.47 16.86 58.56 60.85 61.32 46.66 17.35 52.98 55.74
Add 2 dec layers 80.80 71.65 27.03 67.17 70.51 83.52 75.27 25.48 69.63 72.48 76.57 65.19 29.24 64.04 67.61
Add 2 dec layers 71.44 59.03 18.61 59.88 62.88 75.34 63.06 18.29 62.77 65.55 66.48 52.41 19.78 56.79 59.80
Add 4 dec layers 79.58 70.21 26.45 65.34 69.51 82.27 73.68 25.39 67.82 71.44 74.23 63.78 28.28 62.28 66.02
Add 4 dec layers 71.37 58.79 18.52 59.46 62.54 74.26 63.07 19.12 62.05 65.02 65.48 51.43 19.60 55.62 58.86
End-to-end DIT 81.78 73.07 28.27 68.07 71.46 85.28 77.17 27.56 70.81 73.90 78.02 67.60 30.20 65.58 69.36
End-to-end DIT 76.51 66.00 23.19 62.96 66.87 78.89 68.41 21.31 63.74 67.83 72.90 60.01 25.39 60.43 64.63
Layer-wise DIT 85.68 76.61 29.89 71.98 74.73 88.06 79.17 25.94 74.51 75.62 81.28 69.11 30.95 68.77 71.21
Layer-wise DIT 81.22 70.21 24.31 67.50 69.97 83.89 74.66 24.36 69.55 72.44 77.26 65.44 28.26 64.77 68.12
Table 2. The combination with conventional prompt tuning. Here, dynamic refers to the projected text words changed according to different examples, i.e., the setting of DITs, and static denotes the learnbale tokens.
Setting Prompt val testA testB
P@0.5 P@0.7 oIoU mIoU P@0.5 P@0.7 oIoU mIoU P@0.5 P@0.7 oIoU mIoU
End-to-end dynamic 81.78 73.07 68.07 71.46 85.28 77.17 70.81 73.90 78.02 67.60 65.58 69.36
End-to-end dynamic+static 82.60 73.89 68.53 72.11 85.76 77.76 71.21 74.12 78.41 67.39 65.64 69.38
Layer-wise dynamic 85.68 76.61 71.98 74.73 88.06 79.17 74.51 75.62 81.28 69.11 68.77 71.21
Layer-wise dynamic+static 81.09 71.46 67.31 70.84 83.83 75.30 69.54 72.78 76.84 65.70 63.97 67.98
3.3. Layer-wise Deep Instruction Tuning
According to Wu et al. (Wu et al., 2023b), it is often difficult for additionally inserted tokens to fully interact with the existing feature sequences in the well learned semantic space, especially under the end-to-end manner. In this case, we also propose an effective layer-wise instruction tuning method for SAM, demoted as L-DIT.
Concretely, we project the text tokens onto the sub-semantic space of each Transformer layer of SAM’ encoder, thereby mitigating the modality gap for better tuning. Given the text features $\mathbf{F}_{t}$, we linearly insert them into each layer of the image encoder, i.e., $\mathbf{F}_{t}^{i}$, and the input sequence is defined by
(8) $\mathbf{F}_{m}^{i}=[\mathbf{F}_{t}^{i};\mathbf{F}_{v}^{i}],$
where $i$ denotes the $i$-th layer.
Similar to Eq. 5-7, the final output visual tokens $\mathbf{F}_{v}^{k}$ as well as the projected text features $\mathbf{F}_{t}^{\prime}$ are fed to the lightweight decoder for mask prediction. Compared with E-DIT, this layer-wise scheme can better project text tokens onto the subspaces of SAM’s encoder, thereby improving the efficiency of text instruction tuning.
4. Experiments
To validate the proposed DITs, we conduct extensive experiments on three highly competitive benchmarks of RIS, i.e., RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016), and also compare L-DIT with a bunch of state-of-the-art (SOTA) methods (Zhu et al., 2022; Yang et al., 2022; Liu et al., 2023b, c) in RIS.
4.1. Datasets and Metrics
RefCOCO (Yu et al., 2016) dataset is collected in an interactive two-player game (Kazemzadeh et al., 2014). There are 142,210 referring expressions and 50,000 segmentation masks in 19,994 images collected from COCO (Lin et al., 2014). It is divided into train, val, testA, and testB with 120,624, 10,834, 5,657, and 5,059 samples, respectively.
RefCOCO+ (Yu et al., 2016) contains 141,564 natural language expressions and 49,856 masks in 19,992 images. Compared to RefCOCO, expressions in RefCOCO+ describe more about attributes of the referent, e.g., color, shape, digits, and avoid using words of absolute spatial location such as left and right.
RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016) has 104,560 expressions for 54,822 objects in 26,711 images. The expressions of RefCOCOg contain 8.4 words on average while that of RefCOCO and RefCOCO+ is only 3.6 and 3.5 words. It includes object appearance and location.
Metrics. We use the overall intersection-over-union (oIoU), the mean intersection-over-union (mIoU), and precision at threshold of 0.5, 0.7, and 0.9 as evaluation metrics. The oIoU is measured by the ratio of the total area of intersection to the total area of union across all test samples. The metric favors large objects over small ones. The mIoU is calculated as the average of the IoU values between predictions and the ground truth for all test samples. This measure is applied uniformly to both large and small objects.
Table 3. The impact of freezing the visual and text encoders. Here, ✓means that the encoder is frozen, while ✗is updated.
Visual Text Method val testA testB
P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU
End-to-end 71.83 57.95 18.83 60.76 62.87 75.41 62.49 19.05 63.61 65.67 68.15 52.43 19.54 57.97 60.23
Layer-wise 80.93 69.26 25.00 67.16 69.56 81.64 71.33 23.00 68.24 70.60 73.98 61.01 25.29 62.70 65.59
End-to-end 76.51 66.00 23.19 62.96 66.87 78.89 68.41 21.31 63.74 67.83 72.90 60.01 25.39 60.43 64.63
Layer-wise 81.22 70.21 24.31 67.50 69.97 83.89 74.66 24.36 69.55 72.44 77.26 65.44 28.26 64.77 68.12
End-to-end 81.78 73.07 28.27 68.07 71.46 85.28 77.17 27.56 70.81 73.90 78.02 67.60 30.20 65.58 69.36
Layer-wise 85.68 76.61 29.89 71.98 74.73 88.06 79.17 25.94 74.51 75.62 81.28 69.11 30.95 68.77 71.21
Table 4. The impact of different text word injections for L-DIT. CrossAtt and FFN refer to mutli-head cross attention and feed-forward network (Vaswani et al., 2017), respectively. Linear projection is the setting of L-DIT.
Setting val testA testB
P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU P@0.5 P@0.7 P@0.9 oIoU mIoU
Linear project 85.68 76.61 29.89 71.98 74.73 88.06 79.17 25.94 74.51 75.62 81.28 69.11 30.95 68.77 71.21
CrossAtt 80.49 70.70 25.63 67.06 70.33 83.30 73.45 24.43 69.67 72.01 76.98 65.33 28.08 64.85 68.13
CrossAtt + FFN 82.07 72.22 27.09 68.26 71.42 84.96 76.29 26.25 71.03 73.52 78.34 66.29 29.43 65.68 69.00
Table 5. The impact of different text encoders for L-DIT.
Text encoder val testA testB
oIoU mIoU oIoU mIoU oIoU mIoU
BERT 67.50 69.97 69.55 72.44 64.77 68.12
CLIP 67.19 69.82 68.99 71.60 64.88 67.67
Table 6. The impact of tuning on SAM. Here, means not tuning, and Mix denotes the joint training with point, box and text [^4] Boxes and points are only processed in the default decoder.. The metric used is oIoU.
Setting Train Prompt val testA testB Inference
Default SAM Point 75.84 78.11 75.38 119ms
Default SAM Box 75.32 75.38 76.60 120ms
Default SAM Text Text 60.74 62.95 56.69 123ms
Layer-wise DIT Mix Point 85.16 85.59 84.51 123ms
Layer-wise DIT Mix Box 87.47 87.43 87.74 123ms
Layer-wise DIT Mix Text 69.35 71.04 66.76 123ms
Layer-wise DIT Text Text 71.98 74.51 68.77 123ms
4.2. Experimental Setups
For visual backbone, we use ViT-B/16 initialized from pre-trained weights of SAM. BERT with 12 layers and a hidden dimension of 768 is used as language encoder. We keep the aspect ratio and resize the image to 512 $\times$ 512. Following previous works (Luo et al., 2020a; Zhu et al., 2022), we apply color augmentation, Gaussian blur and random horizontal flipping to augment the data. To binarize the prediction of RIS, we set a threshold of 0.35. The optimizer is Adam (Kingma and Ba, 2014) with a learning rate of 1e-4, which is multiplied by a decay factor of 0.2 at the 30-th and the 35-th, and the batch size is set to 64. The warm-up epoch is configured to 3. The training process spans 40 epochs. More details can refer to our anonymously released project.
Table 7. Comparison with the state-of-the-art methods on three RIS datasets. Here, Scratch refers to only using the RIS examples of these datasets for training, while Pre-trained denotes the use of a large number of visual grounding examples (Krishna et al., 2017) for pre-training. The metric used is oIoU.
Method Visual Backbone Textual Encoder Pretraining Data RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
Trained from scratch:
MCN(CVPR20) (Luo et al., 2020b) Darknet53 bi-GRU 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
EFN(CVPR21) (Feng et al., 2021) ResNet101 bi-GRU 62.76 65.69 59.67 51.50 55.24 43.01 - -
BUSNet(CVPR21) (Yang et al., 2021) DeepLab-R101 Self-Att 63.27 66.41 61.39 51.76 56.87 44.13 - -
CGAN(MM20) (Luo et al., 2020a) DeepLab-R101 bi-GRU 64.86 68.04 62.07 51.03 55.51 44.06 51.01 51.69
ISFP(MM22) (Liu et al., 2022) Darknet53 bi-GRU 65.19 68.49 62.73 52.70 56.77 46.39 52.67 53.00
LST(CVPR21) (Jing et al., 2021) Darknet53 bi-GRU 65.43 67.76 63.08 54.21 58.32 48.02 54.40 54.25
ReSTR(CVPR22)(Kim et al., 2022) ViT-B Transformer 67.22 69.30 64.45 55.78 60.44 48.27 - -
LAVT(CVPR22) (Yang et al., 2022) Swin-B BERT 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
VLT(TPAMI22) (Ding et al., 2022) Swin-B BERT 72.96 75.96 69.60 63.53 68.43 56.92 63.49 66.22
VDP(ICCV23) (Zhao et al., 2023) VQGAN CLIP 73.25 - - 62.69 - - 61.96 -
ReLA(CVPR23) (Liu et al., 2023b) Swin-B BERT 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
SADLR(AAAI23) (Yang et al., 2023) Swin-B BERT 74.24 76.25 70.06 64.28 69.09 55.19 63.60 63.56
CARIS(MM23) (Liu et al., 2023f) Swin-B BERT 75.14 78.10 71.75 66.63 72.04 58.51 64.31 65.82
Pre-trained:
MaLT(arXiv21) (Li et al., 2021) ViLT BERT 3.88M 70.13 71.71 66.92 62.23 65.92 56.06 62.45 62.87
CRIS(CVPR22) (Wang et al., 2022) CLIP-R101 CLIP 3.88M 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
ETRIS(ICCV23) (Xu et al., 2023a) CLIP-R101 CLIP 3.88M 71.06 74.11 66.66 62.23 68.51 52.79 60.28 60.42
SeqTR(ECCV22) (Zhu et al., 2022) Darknet53 bi-GRU 6.10M 71.70 73.31 69.82 63.04 66.73 58.97 64.69 65.74
SEEM(NeurIPS23) (Zou et al., 2023) DaViT-d5 Florence 5.00M - - - - - - 65.70 -
LISA-7B(arXiv23) (Lai et al., 2023) ViT-H LLaVA-7B 18.16M 74.90 79.10 72.30 65.10 70.80 58.10 67.90 70.60
PolyFormer(CVPR23) (Liu et al., 2023c) Swin-L BERT 5.70M 75.96 78.29 73.25 69.33 74.56 61.87 69.20 70.19
Layer-wise DITB (Ours) ViT-B BERT 71.98 74.51 68.77 59.97 65.52 51.72 60.18 61.15
Layer-wise DITH (Ours) ViT-H BERT 76.18 78.13 73.27 68.00 71.77 60.04 67.37 67.92
4.3. Quantitative Analysis
4.3.1. Comparison with the default scheme
In Tab. LABEL:Ablation_of_Tuning_method, we first compare our deep instruction tuning (DIT) methods with the default fusion scheme of SAM and its alternatives, i.e., adding more layers to the mask decoder. We can see that the default fusion scheme performs much worse than DITs on RefCOCO under the setting of either fully tuning or freezing backbone, suggesting that the default lightweight decoder is insufficient in text following. With more decoding layers, this issue can be alleviated to some extent, but the performance still lags behind DITs especially when the image encoder is frozen. To explain, the additional fusion layers can enhance the cross-modal interactions, but also undermines the powerful segmentation capability of SAM pre-trained on massive data. In contrast, without modifying the main structure of SAM, our DIT methods can improve the performance greatly. Another interesting finding from Tab. LABEL:Ablation_of_Tuning_method is that the advantage of layer-wise DIT is more obvious under the setting of freezing backbone. This result confirms our assumption that layer-wise DIT can better help to project text tokens onto the semantic space of SAM. Overall, Tab. LABEL:Ablation_of_Tuning_method well validates the effectiveness of DITs for text-instructed SAM.
4.3.2. Combination with prompt tuning
Tab. LABEL:framwork present the results of combing DITs with conventional the prompt tuning (Zhou et al., 2022a, b; Wu et al., 2023b), i.e., adding learnable tokens to the text prompts. Under end-to-end tuning, the addition of learnable tokens can improve performance to a certain extent, mainly due to the increase of the parameter capacity of DIT for better text tuning. However, these prompt tokens are not that helpful for layer-wise DIT, on the contrary, reducing performance greatly. This result may suggests that the text words are well projected onto each layer of SAM with independent projectors. The combination with parameterized tokens somewhat declines their semantics. These results also indicate the difference between DITs and prompt tuning, i.e., one is for semantic projection while the other is for task adaption.
4.3.3. The impact of different text encoder on Layer-wise DIT
Tab. LABEL:text_encoder examines the impact of different language encoders. It can be seen that BERT outperforms CLIP in most cases, though the gap between the two encoders is not significant. For example, CLIP performs slightly better than BERT in terms of oIoU on testB. We believe this is due to the typically concise descriptions in the dataset, resulting in a relatively limited impact of the text encoder on DIT-SAM.
4.3.4. The effect of freezing encoders
In Tab. LABEL:backbone, we also ablate DITs under the parameter efficient setting (Luo et al., 2023; Wu et al., 2023e). It can be seen that when freezing the SAM’s image backbone and the directly plugged-in BERT, L-DIT can still achieve 67.16% oIoU on RefCOCO val, which is even much better than fully tuning the default SAM, as shown in Tab. LABEL:Ablation_of_Tuning_method. Unfreezing both encoders can greatly improve performance. But compared with the pre-trained BERT, the significance of updating image encoder is more obvious, e.g., +4.88% oIoU on val. This result suggests that in DIT, the image encoder is capable of visual feature learning and cross-modal interactions, well confirming our intuition about DIT.
4.3.5. Different text word injections
In our layer-wise DIT, we project the text prompts directly onto the visual space of each SAM’s layer. Here, we also test DIT with alternative manners in Tab. LABEL:project, such as cross-attention and cross-attention+FFN. Interestingly, these complex methods do not further enhance SAM’s text-aware capability, but decline performance to some extent. This case suggests that with the strong dependency modeling of SAM’s ViT encoder, a direct semantic projection can already facilitate cross-modal alignment, well confirming the motivation of our DIT.
Figure 4. Predictions of the default SAM and our end-to-end and layer-wise DITs, i.e., E-DIT and L-DIT. Both methods can follow text instructions better than the default SAM, while L-DIT can accurately segment the referent even with complex text content or image background. For better comparison, we use red boxes to ground the incorrect segmentations.
4.3.6. The impact of tuning on SAM
In Tab. LABEL:tuning, we further examine the default capability of SAM with point and box guidances on the RefCOCO benchmark. Here, we use the ground-truth boxes of RefCOCO for box-guided SAM, and randomly sample 5 points on the ground-truth mask for point-guided segmentation. From Tab. LABEL:tuning, we can first see that on RefCOCO, the default SAM is still likely to segment the incorrect regions of the referent with ground-truth boxes. To explain, RefCOCO is a challenging benchmark, of which images often involve dense visual semantics, and a large box will make SAM prone segment irrelevant region. In contrast, the point sampled from the object offers a better guidance, but it still lags behind SOTA performance of RIS (Liu et al., 2023c, f). Meanwhile, the joint tuning with all types of prompts can further improve the performance of SAM on following boxes and points. We also notice that our DIT method can help SAM achieve strong text-guided performance that is already close to the box-guided capability, well validating the effectiveness of our method. In terms of the inference speed, since the number of text tokens is often about 15-20, much fewer than that of image patches, the overall impact of DITs on inference time is limited.
4.3.7. Comparison with the State-of-the-art
In Tab. 7, we compare our L-DIT against the state-of-the-art (SOTA) methods of referring image segmentation on RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016) datasets using the oIoU metric. Specially, on RefCOCO dataset, we outperforms existing SOTA methods trained from scratch with absolute improvements of 1.94%, 1.56%, 2.91% in all three splits and performs on par with the pretrained PolyFormer (Liu et al., 2023c). Furthermore, on RefCOCO+ and RefCOCOg datasets, the L-DITH also exhibits exceptional performance, demonstrating comparable capabilities to SOTAs (Liu et al., 2023c; Yang et al., 2023). In addition, compared with SOTA RIS methods using a large number of visual grounding data for pre-training, Our L-DITH is very competitive and achieves new SOTA performance on RefCOCO val.
4.4. Qualitative Analysis
To gain deep insights into DIT, we visualize the predictions of our DIT tuning methods and the default SAM in Fig.4. In the first column, it becomes evident that the default SAM struggles with a nuanced understanding of text, often yielding predictions that inadequately encapsulate the referent or outright incorrect. In contrast, the simple E-DIT offers a notable improvement of SAM’s text comprehension capabilities. For instance, in the third example, SAM’s segmentation exhibits a pronounced bias towards the distracting elements on the left, while E-DIT rectifies this issue, resulting in more accurate prediction. Notably, through E-DIT enhances the performance in some cases, it can lead to the loss of certain semantic information embedded within the textual features. As a consequence, E-DIT occasionally misinterprets the segmentation target, as illustrated in the last example. The last column reveals that L-DIT effectively mitigates these issues without necessitating specific designs for cross-modal interactions. Remarkably, even without providing explicit image positions, akin to the original SAM, L-DIT consistently delivers superior segmentation results. This demonstrates that L-DIT can effectively harness positional information from both modalities and maintain a strong degree of context-awareness, improving accuracy and robustness.
5. Conclusion
In this paper, we propose two deep instruction tuning (DIT) methods to address the weakness of SAM on following text instructions, one is end-to-end and the other is layer-wise. DITs regard SAM’s default visual encoder as a stand-alone multi-modal learner instead of adding additional fusion branches. Both approaches contribute to enhancing SAM’s ability of segmenting target objects based on textual guidance. To validate DITs, we conduct extensive experiments on three RIS benchmark datasets. Experimental results show that our simple DIT methods can greatly improve the capability of SAM. Meanwhile, the proposed layer-wise DIT can even help SAM compete with a set of SOTA methods of RIS on all three datasets.
Acknowledgements.
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).
Appendix A Quantitative Analysis
Figure 5. The attention maps of the default SAM and our end-to-end and layer-wise DITs, i.e., E-DIT and L-DIT. Both methods can follow text instructions better than the default SAM and adjust the area of interest according to text instructions.
A.1. Training Memory Consumption
We compare the memory overhead of L-DIT and existing methods in Tab.  LABEL:memory. From this table, we can see that the additional cost of backbone frozen is close to linear probe tuning and smaller than the full finetuning. In addition, the GPU overhead of full finetuning is even much smaller than LAVT (Yang et al., 2022) in referring image segmentation, since L-DIT does not require another deep fusion model. Compared with LAVT, our L-DIT is simpler and more intuitive, which only requires linear projections to insert text words into each SAM’s layer. These results further validate the efficiency of L-DIT.
Table 8. The memory consumption of L-DIT and existing methods. Here, ✓means that the encoder is frozen, while ✗is updated.
Model Visual Text Batch Size GPU memory
L-DIT 32 11.9G
L-DIT 32 12.4G
L-DIT 32 17.1G
LAVT (Yang et al., 2022) 8 24G
ReLA (Liu et al., 2023a) 32 12G
A.2. The impact of shared parameters for the injection of text features into the visual backbone
In the visual backbone, the shallow and deep layers typically extract low- and high-level features, respectively. To validate this intuition, we experiment on the progressive cascade structure in Tab. LABEL:cascade. From this table, we see that there is no obvious difference between layer-wise and cascade structures. To explain, DIT aims to project the text words onto the visual semantic space of SAM, and simple linear projections may be enough. Similar findings can be also found on recent MLLMs (liu2024visual), where the bridge part is often a stack of linear layers.
Table 9. The impact of shared parameters for the injection of text features into the visual backbone. Cascade means linear layers replace with progressive cascade structure.
Structure val testA testB
mIoU oIoU mIoU oIoU mIoU oIoU
Layer-wise DIT 69.97 67.50 72.44 69.55 68.12 64.77
Cascade DIT 69.94 66.76 72.20 69.72 66.93 63.38
Appendix B Qualitative Analysis
To gain deep insights into DIT, we visualize the attention maps of our DIT tuning methods and the default SAM in Fig. 5. We observe some interesting findings about the self-attention of SAM backbone after inserting text tokens.
For instance, under the frozen backbone setting, the inserted text tokens mainly interact with the visual ones on the higher layers of SAM’s encoder, e.g., the 8-th or 11-th layers. This suggests that the text mainly interacts with the high-level visual semantics of SAM when only tuning the DIT projections.
In contrast, under the full tuning setting, the interactions mainly happen at the shallow layers, e.g., 2-nd and 4-th layers. It might indicate that the text information is early embedded into the visual tokens and then for the deep multi-modal fusion in the image encoder of SAM.
Moreover, it can be observed from the visual attention map that for each layer of the visual backbone, the attentive area of the text instruction is different. For default SAM, it has no obvious interested area, which mainly divides the image into foreground and background. However, for E-DIT and L-DIT, it can be clearly seen that each layer has its attention tendency for different text instructions. The initial layers are interested in entity nouns in text instructions. The 4-th layer is more sensitive to comparative words. At 11-th layer, it mainly focuses on digital information. The 12-th layer is more inclined to color information.